# 1.75cm1.75cm2cm2cm

### v2

@@ 4 Global effects on local 2nd stable pedestals
It becomes apparent from Section(3) that DIII-D pedestals can be close to the 2nd stable region under certain conditions. In fact in various experimental cases, it is observed that plasma parameters do exist where the pedestal is ideal ballooning 2nd stable. Therefore, if both local IBMs and KBMs are found stable an other mechanism should exist that limits the pressure gradient in those cases. ITGs, ETGs and MTMs, if unstable, do not necessarily limit the pressure as they are driving predominantly thermal transport and the pedestal pressure gradient can keep evolving through density built up. TEMs can certainly be found unstable in the pedestal but not enough information exists regarding their impact on pedestal transport. Global gyro-kinetic simulations in JET demonstrated that global effects lead to unstable KBMs in 2nd stability where local KBMs are stabilized. However, performing global gyro-kinetic simulations to examine this effect is impractical due to the immense computational expense. An alternative approach relies on the global ideal MHD model, that is found to provide a destabilizing mechanism through the coupling of kink/peeling modes and finite-$n$ effects on ballooning modes in 2nd stability.
To demonstrate this, a set of DIII-D plasma discharges are considered as test cases and the ELITE code is used to calculate the stability of high-$n$ modes. When $n\sim(\rho^{*}q)^{-1}$, where $\rho^{*}=\rho_{s}/r$ with $\rho_{s}=\sqrt{2m_{s}T_{s}}/(eZ_{s}B)$, the global mode is nearly local in nature as it is localized solely in the steep gradient of the pedestal. For DIII-D pedestals, $k_{y}r/q\sim n=40-60$ depending on the pedestal temperature and $q$-profile ($\rho^{*}$, $r$ and $q$ are taken at the middle of the pedestal). In this analysis, modes with $k_{y}\rho_{s}<1/2$ are assumed to be non-local as their extent can be larger than the pedestal, while modes with $k_{y}\rho_{s}>1/2$ will be stabilized by finite ion Larmor radius effects and are therefore not considered. ELITE is used to find the critical $\beta_{p,ped}$ where $k_{y}\rho_{s}=1/2$ modes are marginally unstable. It is interesting to note that the kink/peeling contribution is necessary for the instability even at such high-$n$, as the ballooning drive alone is not enough to overcome the stabilizing field-line bending.
@@ 5 Application of kinetic and finite-$n$ effects on EPED
To further demonstrate the application of this approach, a number of DIII-D pedestal cases are examined with EPED using GFS, ELITE and BALOO as the KBM constraint. It should be mentioned that in the majority of the pedestals under consideration ideal stability suggests 2nd stability acces which is closed be kinetic or finite-$n$ effects. As it can be observed from Figure(9) better agreement is observed for both pedestal pressure $P_{ped}=2n_{e,ped}T_{e,ped}$ and width $\Delta_{\psi_{N}}$ between EPED1 (EXP/ELITE), EPED1.63 (BALOO/ELITE), EPED2 (ELITE/ELITE) and EPED3 (GFS/ELITE) with experimental measurements. In particular, the standard deviation of EPED1 is $\sigma_{P}=0.14\%$ and $\sigma_{\Delta}=0.21\%$, for EPED1.63 is $\sigma_{P}=0.17\%$ and $\sigma_{\Delta}=0.26\%$, for EPED2 is $\sigma_{P}=0.12\%$ and $\sigma_{\Delta}=0.2\%$, while for EPED3 is $\sigma_{P}=0.08\%$ and $\sigma_{\Delta}=0.15\%$. It can be concluded that using GFS or ELITE to compute the KBM constraint led to better agreement with experimental measurements and uncertainty for both pedestal pressure and width, highlighting the importance of kinetic or finite-$n$ effects close to 2nd stability.
@@ 6 Summary
This work focused on precisely and efficiently calculating the KBM pedestal width $\Delta_{\psi_{N}}$ and height $\beta_{p,ped}$ considering that typically takes the form of $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$. The required improvements are motivated by the unique characteristics of NSTX(-U) pedestals, that are typically not described by the IBM stability, and the recent resolution of the discrepancy attributed to kinetic effects. The development of the GFS code allowed for an accurate and fast numerical algorithm to perform KBM stability calculations as it is shown through this work.
Experimental relevant pedestal width/height scaling is reproduced considering the KBM critical $\beta_{p,ped}$ as calculated from GFS for a number of DIII-D and NSTX-U plasmas. A comparison between DIII-D and NSTX-U showed that the DIII-D pedestal are found close to the nose of the $\hat{s}-\alpha$ stability diagram, while NSTX-U pedestals are found at higher $\hat{s}$ in the 1st stability boundary for the cases under consideration. Through changes on the plasma parameters and geometry of NSTX-U #139047, the originally calculated linear width/height scaling approached the square root dependence when the pedestal approached the nose of the stability diagram.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, Theory Program and the NSTX-U Research Program, using the NSTX-U Fusion Facility, a DOE Office of Science user facility. We thank the NSTX-U experimental team for providing the data analyses.

### v3

@@ 4 Global effects on local 2nd stable pedestals
It becomes apparent from Section(3) that DIII-D pedestals can access the 2nd stable region under certain conditions. In fact in various experimental cases, it is observed that plasma parameters do exist where the pedestal is ideal ballooning 2nd stable. Therefore, if both local IBMs and KBMs are found stable another mechanism should exist that limits the pressure gradient in those cases. ITGs, ETGs and MTMs, if unstable, do not necessarily limit the pressure as they are driving predominantly thermal transport and the pedestal pressure gradient can keep evolving through density built up. TEMs can certainly be found unstable in the pedestal but not enough information exists on their impact as a pedestal constraint. Global gyro-kinetic simulations in JET demonstrated that global effects lead to unstable KBMs in 2nd stability where local KBMs are stabilized [11]. However, performing global gyro-kinetic simulations to examine this effect is impractical due to the immense computational expense. An alternative approach relies on the global ideal MHD model, that is found to provide a destabilizing mechanism through the coupling of kink/peeling modes and finite-$n$ effects on ballooning modes in 2nd stability [12][13].
To demonstrate this, a set of DIII-D plasma discharges are considered as test cases and the ELITE code is used to calculate the stability of high-$n$ modes. When $n\sim(\rho^{*}q)^{-1}$, where $\rho^{*}=\rho_{s}/r$, the global mode is nearly local in nature as it is localized solely in the steep gradient of the pedestal. In this analysis, modes with $k_{y}\rho_{s}<1/2$ are assumed to be non-local as their extent can be larger than the pedestal, while modes with $k_{y}\rho_{s}>1/2$ will be stabilized by finite ion Larmor radius effects and are therefore not considered. ELITE is used to find the critical $\beta_{p,ped}$ where $k_{y}\rho_{s}=1/2$ modes are marginally unstable. It is interesting to note that the kink/peeling contribution is necessary for the instability even at such high-$n$, as the ballooning drive alone is not enough to overcome the stabilizing field-line bending. For typical DIII-D pedestals, it is found that $n\sim 40-60$ depending on the pedestal temperature and $q$-profile ($\rho^{*}$, $r$ and $q$ are taken at the middle of the pedestal) satisfy $k_{y}\rho_{s}=1/2$.
@@ 5 Application of kinetic and finite-$n$ effects on EPED
To further demonstrate the application of this approach, a number of DIII-D pedestal cases are examined with EPED using GFS, ELITE and BALOO as the KBM constraint. It should be mentioned that in the majority of the pedestals under consideration, ideal stability suggests that significant part of the pedestal has 2nd stability access which is limited by kinetic or finite-$n$ effects. As a result,t the BALOO calculation might not provide an accurate scaling. For each plasma pulse GFS, ELITE and BALOO are used to evaluate the pedestal width/height scaling, which is then used in EPED to compute the global MHD boundary with ELITE. The standard scaling of EPED1, where $\Delta_{\psi_{N}}=0.076\beta_{p,ped}^{1/2}$, is also compared. As it can be observed from Figure(9) better agreement is observed for both pedestal pressure $P_{ped}=2n_{e,ped}T_{e,ped}$ and width $\Delta_{\psi_{N}}$ between EPED1 (EXP/ELITE), EPED1.63 (BALOO/ELITE), EPED2 (ELITE/ELITE) and EPED3 (GFS/ELITE) with experimental measurements when using the global or kinetic constraint width/height. In particular, the standard deviation of EPED1 is $\sigma_{P}=0.14\%$ and $\sigma_{\Delta}=0.21\%$, for EPED1.63 is $\sigma_{P}=0.17\%$ and $\sigma_{\Delta}=0.26\%$, for EPED2 is $\sigma_{P}=0.12\%$ and $\sigma_{\Delta}=0.2\%$, while for EPED3 is $\sigma_{P}=0.08\%$ and $\sigma_{\Delta}=0.15\%$. It can be concluded that using GFS or ELITE to compute the KBM constraint led to better agreement with experimental measurements for both pedestal pressure and width, highlighting the importance of kinetic or finite-$n$ effects close to 2nd stability.
@@ 6 Summary
This work focused on calculating the KBM pedestal width $\Delta_{\psi_{N}}$ and height $\beta_{p,ped}$ considering that typically takes the form of $\Delta_{\psi_{N}}=c_{1}\beta_{p,ped}^{c_{2}}$. The required improvements are motivated by the unique characteristics of NSTX(-U) pedestals, that are typically not described by the IBM stability, and the recent resolution of the discrepancy attributed to kinetic effects. The development of the GFS code allowed for an accurate and fast numerical algorithm to perform KBM stability calculations as it is shown through this work.
Experimentally relevant pedestal width/height scaling is reproduced considering the KBM critical $\beta_{p,ped}$ as calculated from GFS for a number of DIII-D and NSTX plasmas. A comparison between DIII-D and NSTX showed that the DIII-D pedestals are close to the nose of the $\hat{s}-\alpha$ stability diagram, while NSTX pedestals are found at higher $\hat{s}$ in the 1st stability boundary for the cases under consideration. Through changes on the plasma parameters and geometry of NSTX #139047, the originally calculated linear width/height scaling approached the square root dependence when the pedestal approached the nose of the stability diagram.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, Theory Program and the NSTX Research Program, using the NSTX Fusion Facility, a DOE Office of Science user facility. We thank the NSTX experimental team for providing the data analyses.
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.